Two two-digit numbers are made from four of the digits 1,2,3,4 and 5 with no digit being used more than once. What numbers would make:
a) the smallest product?
b) the largest product?
c) a product closest to 700?
please help me!

Answer by Edwin McCravy(20065)

13×24=312 = the smallest possible product.
43×52=2236 = the largest possible product.
13×54=702 = the closest possible product to 700

Edwin